OpenStudy (anonymous):

2sinx-3cosx/4sinx-9cos when cosecx = 13/12

OpenStudy (ajprincess):

 cosecx=hyp/opp \(144+{adj}^2=169\) \({adj}^2=169-144\) \({adj}^2=25\) \(adj=5\) \[sinx=\frac{opp}{hyp}\] \[sinx=\frac{12}{13}\] \[cosx=\frac{adj}{hyp}\] \[cosx=\frac{5}{13}\] \[\frac{2sinx-3cosx}{4sinx-9cos}\] \[=\frac{2*\frac{12}{13}-3*\frac{5}{13}}{4*\frac{12}{13}-9*\frac{5}{13}}\]